OpenStudy (anonymous):

A waiter earned $31 in tips in an hour. He was paid in $1 bills and $5 bills. Use a model to determine how many $5 bills he received if he received a total of 15 bills.

OpenStudy (campbell_st):

well let F = the number of five dollar bills O = the number of one dollar bills the total number of bills is 15 so F + O= 15 now for the value of the bills is $31 then 5F +O = 31 you have 2 equations in 2 unknowns that you now need to solve to find F

OpenStudy (anonymous):

let a = the number of 1 dollar bills, let b = the number of five dollar bills. so \[31 = a+5b\] and \[a + b = 15\]

OpenStudy (anonymous):

solve the second equation for a and then plug the result into the first equation

OpenStudy (campbell_st):

exactly the same... you can solve by substitution or elimination

OpenStudy (anonymous):

yup, the same :)

OpenStudy (campbell_st):

yep... thats substitution method and that works

OpenStudy (anonymous):

i dont get it can you break it down a little..please

OpenStudy (anonymous):

"total of 15 bills " -> O + F = 15

OpenStudy (campbell_st):

well you have 15 notes made up of $5 and $ 1 so you say F + O = 15 does that make sense

OpenStudy (anonymous):

@kardaja Hello...!!

OpenStudy (campbell_st):

you also need to set up an equation for the value of the notes so 5F + O = 31 or 5 times the number of five dollar bills + 1 times the number of one dollar bills has a total value of $31.

OpenStudy (anonymous):

Okay. Thank You.

OpenStudy (anonymous):

No of $5 bills No of $1 bills 1 26 NO 2 21 NO 3 16 NO 4 11 YES! 5 6 No need to do more (This only works because it's small set)

OpenStudy (anonymous):

Let x= one dollar bills and y= $5 bills x+y=15 x+5y=31 4y=16 y=4 ($5 bills)=$20 x=11 ($1 bills)= 11 15bills $31

OpenStudy (anonymous):

he has 4 bills